Is 2 a solution of 14x - 16 ?
Answer by mixerimaythyme(7)

2 is not a solution to 14x-16=0
since it is a linear equation, there is only one solution for x
take 14x-16=0 and add 16 to both sides, which gives you 14x=16
now divide 14 on both sides, which gives you x= (16)/ (14) which siplifies to...
x= 8/7
Okay?